# Cognitive Foundations for Reasoning and Their Manifestation in LLMs

### v1

@@ 3.2.1 Dataset Composition
The distribution of problem types (Figure 2(a)) varies substantially across modalities but decrease as the problem becomes less structured. Algorithmic problems dominate with  6,300 text, 8,400 image, and 800 audio samples. Among structured types, Rule-Using shows strong concentration in images (4,447 samples) versus text (521) and audio (248). Notably, Strategic Performance has zero instances across all modalities, while the LLM classifiers frequently considered this category with explanations including "I considered Strategic Performance but rejected it because this is an offline planning problem under perfect information, not real-time execution." Strategic Performance necessitates interactive benchmarking, which is not feasible under our experimental constraint to generate reasoning traces to the same input across models, but the lack of this problem type in the single-turn, static benchmarks echos the call for interactive evaluation of models (hofmann2025fluid; li2024mediq). Three problem types fall outside the structuredness ranking: Factual Recall is the most prevalent, showing balanced representation across modalities (2,841 text, 2,663 image, 3,307 audio) and representing the second-largest category overall. Logical problems, involving abstract reasoning, are heavily image-focused (2,215 samples) but rare in text (57) and audio (152). Overall, text reasoning shows the broadest distribution across problem types, while image and audio concentrate on Algorithmic, Rule-Using, Logical, and Factual Recall.
@@
Figure 3: Dataset composition and model performance across problem types. (a) Problem distribution across modalities, organized by Jonassen’s problem structuredness continuum, shows coverage decreases for less structured problems. (b) Accuracy decreases as problems become less structured, with models showing consistent performance on algorithmic tasks (63.8%) but high variance on dilemmas (3.3-99.1%).
@@
Success rates vary substantially by problem structure (Figure 2(b)). Well-structured problems achieve higher correctness overall (algorithms: 63.8%, story problems: 79.5%), while ill-structured problems average lower success (design: 46.6%). Dilemma tasks show polarization (3.2-99.1%) depending on the model size. Detailed model-specific performance breakdowns appear in Appendix A.3.
@@ 3.2.2 Analysis Dimensions
To thoroughly examine the manifestation of behaviors within reasoning traces, we organize our analysis around three interconnected research questions:
@@
    Which behaviors are most prevalent in reasoning traces, and how does their frequency relate to reasoning success? We investigate the distribution of reasoning behaviors across different models and problem types, examining whether prevalence patterns vary along Jonassen’s well-structured to ill-structured problem continuum. Critically, we distinguish between behaviors that are frequently exhibited and those that are most strongly correlated with correct outcomes, as these sets may not align. This analysis reveals whether models consistently employ the behaviors most conducive to success or instead are biased towards alternative strategies.
@@
    What structural dependencies exist between reasoning behaviors? Beyond examining individual behavior presence, we analyze their temporal and hierarchical interdependencies within reasoning traces. The mere presence of success-correlated behaviors does not guarantee effective reasoning; rather, the ordering and composition of behaviors critically influences problem-solving efficacy. For instance, when resolving a dilemma, failure to employ heightconceptual processing to heightdecompose the problem into constituent considerations early in the reasoning process may fundamentally compromise the efficiency and quality of the eventual judgment, even if these behaviors appear later. To identify optimal behavior sequences, we leverage our transition graph representation (Section 3.1.4) to extract consensus subgraphs—common structural patterns that maximize collective NPMI scores across successful traces, thereby revealing the behavioral scaffolding most strongly associated with correct reasoning.
@@
    How do behavioral structures differ between LLMs and human reasoners? We conduct a comparative study examining both the distributional characteristics of behaviors exhibited by LLMs versus humans on a shared problem set, and perform fine-grained, span-level qualitative analysis to understand how they differ in their reasoning processes and behavior utilization strategies.
@@ 3.3.1 Distribution of Cognitive Elements
The first phase of our analysis investigates the misalignment between cognitive elements that models frequently deploy (through behavioral manifestation) versus those that correlate most strongly with success. Figure 4 presents this contrast: the left heatmap shows behavioral prevalence across all traces for each problem type, while the right displays the positive pointwise mutual information (PPMI) between behavior occurrence and trace success.
@@
Models deploy behaviors inversely to what success requires. The contrasting patterns between the two heatmaps reveal a fundamental misalignment in how models adapt their reasoning strategies. On well-structured problems (e.g., algorithmic, rule-using), models deploy a broad repertoire of behaviors at high frequencies, where the average presence across all behaviors is $0.3652\pm 0.2598$ for algorithmic, story, and rule-using problems. As problems become ill-structured and non-verifiable, however, models narrow their behavioral repertoire: ill-structured problems like case analysis, design, and dilemma have a $0.2787\pm 0.2672$ average presence across all behaviors, with usage concentrating heavily on heightsequential organization, heightlogical coherence, and heightforward chaining.
@@
Figure 4: (Left) Presence rate of each reasoning behavior for each problem type (ranging from well-structured to ill-structured). (Right) Positive Pointwise Mutual Information (PPMI) between the problem type and reasoning behavior (correlation between the behavior occurrence and reasoning trace success).
@@
This narrowing strategy is the opposite of what is reflected in successful traces. The right heatmap reveals that overall success on ill-structured problems demands greater behavioral diversity. As we move from well-structured to ill-structured problem types, the PPMI scores become more uniformly elevated across a wider range of behaviors, particularly diverse heightrepresentations (hierarchical, network, spatial, temporal) and varied heightoperations (backward chaining, representational restructuring, pattern recognition). Specifically, algorithmic, story, and rule-using problems have an average PPMI score of $0.052\pm 0.067$ across all behaviors, while the last four ill-structured problems (from diagnosis-solution to dilemma) have an average score of $0.224\pm 0.182$.
Thus, while well-structured problems, which are more tolerant of uniform approaches, receive broad behavioral engagement, ill-structured problems, which critically require a diverse repertoire of reasoning behaviors, only receive a subset dominated by heightsequential processing and heightforward chaining. This inverse relationship between behavioral deployment and success reflects that models have learned to apply their most diverse reasoning behaviors where they are the least necessary, while defaulting to limited, inflexible strategies precisely where adaptability matters most.
@@
Models frequently attempt core reasoning behaviors but struggle to execute them effectively. Reasoning invariants—particularly heightlogical coherence and heightcompositionality—appear ubiquitously across problem types, yet show surprisingly weak correlation with success. For example, logical coherence’s average NPMI across all problem types is $0.0984$, whereas knowledge alignment features an average score of $0.301$. Manual inspection of traces by human annotators reveals a systematic pattern. While models frequently attempt to identify logical inconsistencies and contradictions, they consistently fail to recognize or effectively respond to them, unlike human reasoners. This execution gap explains the discrepancy between high prevalence and low predictive value of these foundational behaviors.
@@
Models show limited meta-cognitive success, especially on problems lacking clear ground truth. The heightevaluation element similarly demonstrates high prevalence but negligible success correlation, with its frequency declining sharply for ill-structured problems (case analysis, dilemma). This pattern suggests models struggle particularly with self-assessment on non-verifiable problems where ground truth is ambiguous or absent.
@@
Default sequential organization hampers performance when problems require alternative representations. Models exhibit strong preference for heightsequential and heightcausal organization regardless of problem type. However, the success patterns indicate that as the problem type becomes less structured (algorithmic $\rightarrow$ dilemma), diverse organizational strategies—including hierarchical, network, ordinal, temporal, and spatial representations—become increasingly critical. This reflects a fundamental challenge of ill-defined problems: when problem descriptions lack inherent structure, successful reasoners must actively construct appropriate organizational frameworks rather than defaulting to sequential processing.
@@
Operational rigidity limits models’ ability to adapt reasoning strategies to problem demands. Similar rigidity emerges in reasoning operations. Models consistently favor heightselective attention, heightdecomposition and integration, and heightforward chaining, the latter being a natural consequence of their sequential organizational bias. Successful traces, however, demonstrate substantially greater operational diversity, adapting their reasoning strategies to problem characteristics rather than applying uniform approaches.
@@
Model-Specific Behavior Distribution
To complement our problem-type analysis, we examine how behavioral deployment varies across different model architectures and modalities. Figure 6 displays the presence rate of each reasoning behavior across eleven models, including text-only systems (the majority), audio-capable models (Qwen3-Omni), and vision-language models (Zebra-CoT).
Several behaviors exhibit consistently high presence rates across nearly all models, suggesting core reasoning primitives that emerge regardless of architecture or training methodology. heightBacktracking] appears frequently across all models (presence rates exceeding 0.8 across all models), indicating that iterative refinement and error correction are prominent components of model reasoning processes.
With respect to the different modalities, we observe that the audio and image models have overall less presence of reasoning behaviors, exhibiting certain behaviors frequently (e.g., heightevaluation, heightproductivity) and others minimally, relative to the textual models. This indicates that there may be lower diversity of reasoning behaviors present in multimodal models.
Figure 5: The most successful vs. most common 7-node behavioral pattern across all Diagnosis-Solution reasoning trace with the correlation to success (NPMI) and the probability of each node.
3.3.2 Behavioral Structures
Our analysis in Section 3.3.1 demonstrated that successful reasoning traces exhibit diverse behaviors, adapting their selection to problem-specific demands. However, behavioral presence alone provides an incomplete picture. Thus, we ask: does the temporal and hierarchical organization of these behaviors also differ between successful and unsuccessful traces? To address this question, we extract optimal behavioral structures for each problem type by analyzing the transition graphs constructed from all model traces $T$.
Behavioral consensus subgraph extraction.
We leverage the graph construction methodology detailed in Section 3.1.4 to encode and aggregate reasoning traces for each problem type, yielding a complete transition graph $G$ where each node represents a behavior and each edge captures their relationship (hierarchical, sequential, or parallel). Each node and edge is weighted by its NPMI score, quantifying the strength of association between that behavioral element and trace success.
Given graph $G$, we extract a consensus subgraph $G^{*}\subseteq G$ that represents the prototypical successful reasoning structure through a greedy forward search procedure. First, we identify the starting behavior by selecting the node with highest NPMI among behaviors that initiate successful traces (i.e., behaviors with the earliest temporal spans). We denote this initial node as $b_{\text{start}}$. From the current behavior $b_{\text{curr}}$, we iteratively expand the consensus structure by selecting the outgoing edge $(b_{\text{curr}},b_{\text{next}},e_{\text{type}})$ with the maximum NPMI score, where $e_{\text{type}}\in\{\texttt{CONTAINS},\texttt{NEXT},\texttt{PAR}\}$. To ensure acyclicity and prevent redundant paths, we maintain a visited set and exclude any target behaviors already present in $G^{*}$. This process continues until either: (1) we reach a predefined maximum number of nodes $|V_{\max}|$, (2) no valid outgoing edges exist from $b_{\text{curr}}$, or (3) when operating in dynamic mode, all remaining edges have non-positive NPMI scores, indicating they do not positively correlate with success.
@@ 3.3.2 Behavioral Structures
Figure 6: Presence rate of each reasoning behavior for each model (how often is the behavior occurring across all reasoning traces for a model), including all modalities (Qwen3-Omni: audio; Zebra-CoT: image).
The resulting consensus subgraph $G^{*}$ represents the behavioral scaffolding most strongly associated with successful reasoning for the given problem type. Critically, this structure captures not merely which behaviors appear, but how they are sequenced and composed—revealing the temporal dependencies and hierarchical decompositions that characterize effective problem-solving strategies. By comparing $G^{*}$ across problem types, we can identify whether optimal reasoning structures remain consistent or adapt fundamentally to different levels of problem structure.
Analysis.
Figure 5 illustrates the consensus subgraph constructed for Algorithmic problems. We observe that successful traces begin with selective attention to identify relevant problem features, followed by sequential organization and forward chaining to execute solution steps. This structure contrasts sharply with the most common pattern, which prioritizes logical coherence and compositionality early but achieves lower success correlation. The divergence suggests that immediate focus on abstract reasoning principles without first attending to problem-specific constraints leads to less effective problem-solving on algorithmic tasks.

### v2

@@ 3.2.1 Dataset Composition
The distribution of problem types (Figure 2(a)) varies substantially across modalities, decreasing as problems become less structured. Algorithmic problems dominate ( 6,300 text, 8,400 image, 800 audio), while Rule-Using shows strong image concentration (4,447 samples) versus text (521) and audio (248). Strategic Performance has zero instances across all modalities. Since it requires interactive benchmarking infeasible under our experimental design to generate reasoning traces to identical inputs across models, echoing calls for interactive evaluation (hofmann2025fluid; li2024mediq). Three problem types fall outside the structuredness ranking: Factual Recall is most prevalent with balanced cross-modal representation (2,841 text, 2,663 image, 3,307 audio); Logical problems concentrate in images (2,215 samples) but are rare in text (57) and audio (152); Creative/Expressive tasks appear minimally (7 text samples only). Overall, text reasoning shows the broadest problem type distribution, while image and audio concentrate on Algorithmic, Rule-Using, Logical, and Factual Recall.
@@
Figure 3: Dataset composition and model performance across problem types. (a) Problem distribution across modalities, organized by Jonassen’s problem structuredness continuum, shows coverage decreases for less structured problems. (b) Accuracy decreases as problems become less structured, with models showing consistent performance on story tasks (78.8%) but high variance on dilemmas (3.3-99.1%).
@@
Success rates vary substantially by problem structure (Figure 2(b)). Well-structured problems achieve higher correctness overall (algorithms: 62.2%, story problems: 78.8%), while ill-structured problems average lower success (design: 48.0%). Dilemma tasks show polarization (3.2-99.1%) depending on the model size. Detailed model-specific performance breakdowns appear in Appendix A.3.
@@ 3.2.2 Analysis Dimensions
To thoroughly examine the manifestation of cognitive elements within reasoning traces, we organize our analysis around three interconnected research questions:
@@
    Which cognitive elements are most prevalent in reasoning traces, and how does their frequency relate to reasoning success? We investigate the distribution of elements across different models and problem types, examining whether prevalence patterns vary along Jonassen’s well-structured to ill-structured problem continuum. By differentiate between elements that are frequently exhibited and those that are most strongly correlated with correct outcomes, this analysis reveals whether models consistently employ the elements most conducive to success, or are biased towards alternative strategies.
@@
    What structural dependencies exist between cognitive elements? Beyond examining individual behavioral presence, we analyze their temporal and hierarchical interdependencies within reasoning traces. The mere presence of success-correlated elements does not guarantee effective reasoning; rather, the ordering and composition of elements critically influences problem-solving efficacy. For instance, when resolving a dilemma, failure to employ heightconceptual processing to heightdecompose the problem into constituent considerations early in the reasoning process may fundamentally compromise the efficiency and quality of the answer, even if these elements appear later. To identify optimal behavioral sequences, we construct reasoning structure representation (Section 3.1.4) by extracting common structural patterns that maximize collective NPMI scores across successful traces to reveal the behavioral scaffolding most strongly associated with correct reasoning.
@@
    How do reasoning structures differ between LLMs and human reasoners? We conduct a comparative study examining both the distributional characteristics of cognitive elements exhibited by LLMs versus humans on a shared problem set, and perform fine-grained, span-level qualitative analysis to understand how they differ in their reasoning processes and behavioral utilization strategies.
@@ 3.3.1 Distribution of Cognitive Elements
We first investigate the misalignment between cognitive elements that models frequently deploy (through behavioral manifestation) versus those that correlate most strongly with success. Figure 4 shows behavioral prevalence across all traces for each problem type (left) versus the positive pointwise mutual information (PPMI) between behavioral occurrence and trace success (right).
@@
Models deploy elements inversely to what success requires. The contrasting patterns between the two heatmaps reveal a fundamental misalignment in how models adapt their reasoning strategies. On well-structured problems (e.g., algorithmic, rule-using), models deploy a broad repertoire of behaviors at high frequencies, where the average presence across all elements is $0.397\pm 0.255$ for algorithmic, story, and rule-using problems. As problems become ill-structured and non-verifiable, however, models narrow their behavioral repertoire: ill-structured problems like case analysis, design, and dilemma have a $0.337\pm 0.261$ average presence across all elements, with usage concentrating heavily on heightsequential organization, heightlogical coherence, and heightforward chaining.
@@
Figure 4: (Left) Presence rate of each cognitive element for each problem type (ranging from well-structured to ill-structured). (Right) Positive Pointwise Mutual Information (PPMI) between the problem type and cognitive element (correlation between their behavioral occurrence and reasoning trace success).
@@
This narrowing strategy is the opposite of what is reflected in successful traces. The right heatmap reveals that overall success on ill-structured problems demands greater behavioral diversity. As we move from well-structured to ill-structured problem types, the PPMI scores become more uniformly elevated across a wider range of cognitive elements, particularly diverse heightrepresentations (hierarchical, network, spatial, temporal) and varied heightoperations (backward chaining, representational restructuring, pattern recognition). Specifically, algorithmic, story, and rule-using problems have an average PPMI score of $0.046\pm 0.063$ across all elements, while the last four ill-structured problems (from diagnosis-solution to dilemma) have an average score of $0.186\pm 0.114$. Thus, well-structured problems, which are more tolerant of uniform approaches, receive broad behavioral engagement, while ill-structured problems, which critically require a diverse repertoire of cognitive elements, only receive a subset dominated by heightsequential processing and heightforward chaining. This inverse relationship between behavioral deployment and success reflects that models have learned to apply their most diverse cognitive elements where they are the least necessary, while defaulting to limited, inflexible strategies precisely where adaptability matters most.
@@
Models frequently attempt core cognitive elements, but struggle to execute them effectively. Reasoning invariants—particularly heightlogical coherence and heightcompositionality—appear ubiquitously across problem types, yet show surprisingly weak correlation with success. For example, the average PPMI of heightlogical coherence is $0.091$, despite having a presence rate of $91\%$. On the other hand, heightknowledge alignment features an average PPMI of $0.234$ while only being featured in $20.2\%$ of traces. Manual inspection of traces by human annotators reveals a systematic pattern. While models frequently attempt to identify logical inconsistencies and contradictions, they consistently fail to recognize or effectively respond to them, unlike human reasoners. This execution gap explains the discrepancy between high prevalence and low predictive value of these foundational elements.
@@
Models show limited meta-cognitive success, especially on problems lacking clear ground truth. heightEvaluation similarly demonstrates high prevalence ($53.5\%$) but low success correlation (a PPMI of $0.031$), with its frequency declining sharply for ill-structured problems (case analysis, dilemma). This pattern suggests models struggle particularly with self-assessment on non-verifiable problems where the ground truth is ambiguous or absent.
@@
Default sequential organization hinders performance when problems require alternative representations. Models exhibit strong preference for heightsequential and heightcausal organization regardless of problem type. However, the success patterns indicate that as the problem type becomes less structured (algorithmic $\rightarrow$ dilemma), diverse organizational strategies (e.g. hierarchical, network, ordinal, temporal, and spatial representations) become increasingly critical. For example, heightspatial organization has an average presence rate of $9.8\%$, despite a PPMI of $0.252$. This reflects a fundamental challenge of ill-defined problems: when problem descriptions lack inherent structure, successful reasoners must actively construct appropriate organizational frameworks rather than defaulting to sequential processing.
@@
Operational rigidity limits adaptation of reasoning strategies to problem demands. Similar rigidity emerges in reasoning operations. Models consistently favor heightselective attention, heightdecomposition height& integration, and heightforward chaining, the latter being a natural consequence of their sequential organizational bias. Successful traces, however, demonstrate substantially greater operational diversity, adapting their reasoning strategies to problem characteristics rather than applying uniform approaches.
@@
Figure 5: (Left) Presence rate of each cognitive element for each model (how often is the element occurring across all reasoning traces for a model) across all modalities. Average rates per model: DeepSeek-R1: 0.458, Olmo-3-7B-Think: 0.491, Olmo-3-32B-Think: 0.484, Qwen3-8B: 0.357, Qwen3-14B: 0.35, Qwen3-32B: 0.384, DeepSeek-R1-Distill-Qwen-1.5B: 0.316, DeepSeek-R1-Distill-Qwen-7B: 0.334, DeepSeek-R1-Distill-Qwen-14B: 0.349, DeepSeek-R1-Distill-Qwen-32B: 0.36, DeepSeek-R1-Distill-Llama-8B: 0.315, DeepSeek-R1-Distill-Llama-70B: 0.346, DeepScaleR-1.5B-Preview: 0.317, DeepHermes-3-Llama-3-8B-Preview: 0.122, OpenThinker-32B: 0.505, s1.1-32B: 0.597, Qwen3-Omni-30B (audio): 0.253, and Zebra-CoT (image): 0.348. (Right) Average reasoning trace length (# characters) for each model per problem type.
Model-Specific Behavior Distribution.
To complement our problem-type analysis, we examine how behavioral deployment of cognitive elements varies across different model architectures and modalities. Figure 5 (left) displays the behavioral presence rate of each cognitive element across 18 models, including text-only systems (the majority), audio-capable models (Qwen3-Omni), and vision-language traces (Zebra-CoT).
Several elements exhibit consistently high presence rates across nearly all models, suggesting core reasoning primitives that emerge regardless of architecture or training methodology. For example, both heightsequential organization and heightforward chaining appear frequently across all models, indicating that bias towards autoregressive, next-token training paradigms are exhibited in element preferences. Comparing two heatmaps in Figure 5, we observe that there exists a correlation between model-specific behavioral diversity and trace length. Specifically, we note that the models with the highest average presence rate across all behaviors are Olmo-3-7B-Think ($48.4\%$ presence rate; $17,416$ average characters) and s1.1-32B ($59.7\%$ presence rate; $13,210$ average characters). Additionally, we observe that well-structured problem types feature longer reasoning traces. We hypothesize that this may either be due to (a) a well-defined problem may require more reasoning surrounding the provided knowledge and constraints of the problem, or (b) an ill-defined problem has fewer constraints, allowing the model more freedom and leniency to formulate a solution. We note that more robust evaluation mechanisms for open-ended problems will be beneficial to incentivize more thorough, well-reasoned solutions (as there is a strong bias towards verifiable problems in LLM training data and evaluation). Across modalities, the audio and image models have overall less presence of cognitive elements, exhibiting certain elements frequently (e.g., heightevaluation, heightproductivity) and others minimally, relative to the textual models. This indicates that there may be lower diversity of reasoning present in multimodal models.
3.3.2 Reasoning Structures
Our analysis in Section 3.3.1 demonstrated that successful reasoning traces exhibit diverse cognitive elements, adapting their selection to problem-specific demands. However, behavioral presence alone provides an incomplete picture. Thus, we ask: does the temporal and hierarchical organization of these behavioral manifestations also differ between successful and unsuccessful traces? To address this question, we extract the most common and most successful reasoning structures for each problem type by analyzing the transition graphs constructed from all model traces $T$.
We leverage the graph construction methodology detailed in Section 3.1.4 to extract two representative structures for each problem type: $\hat{G}$ representing the prototypical reasoning structure models commonly employ, and $G^{*}$ representing behavioral patterns characteristic of successful traces. Critically, these structures capture not merely which elements appear, but how they are sequenced and composed, highlighting the temporal dependencies and hierarchical decompositions that characterize effective problem-solving strategies. By comparing $\hat{G}$ and $G^{*}$ across problem types, we can identify whether models deploy reasoning structures aligned with success patterns, or rely on different behavioral organizations that may be less effective.
Figure 6 illustrates successful versus common reasoning structures for Algorithmic and Diagnostic problems, revealing systematic misalignment. For Algorithmic problems (well-structured), the most common pattern includes elements with negative NPMI scores: heightself-awareness (-0.141) and heightbacktracking (-0.050), indicating these frequently deployed behaviors actually correlate with failure. Successful traces instead begin with heightselective attention followed by heightlogical coherence and heightsequential organization. For Diagnostic problems (ill-structured), the structural divergence is more pronounced. Successful traces follow a deliberate scoping strategy: heightselective attention $\rightarrow$ heightsequential organization $\rightarrow$ heightknowledge alignment before engaging heightforward chaining. This structure first identifies relevant features and aligns with domain constraints before solution construction. The common pattern bypasses this scoping phase entirely, immediately rushing into heightforward chaining (Prob: 0.748). This premature solution-seeking explains systematic failures on problems requiring constraint satisfaction, in which models generate solutions before understanding what makes them valid. For both problem types, successful reasoning exhibit more diverse structural relationships among the cognitive elements whereas the most common traces exclusively connect the elements sequentially.
Figure 6: Successful vs. Common 7-node behavioral pattern across all Algorithmic and Diagnosis-Solution reasoning trace with the correlation to success (NPMI) and the probability of each node.